OpenStudy (anonymous):

Find the volume of a cylinder with a height of 4 meters and a diameter that is 3 times the measure of the height. Use 3.14 for pi, and round your answer to the nearest hundredth.

OpenStudy (anonymous):

Use the formula V=1/3 3.14r^2 H

OpenStudy (anonymous):

\[ V =\pi r^2h \]

OpenStudy (anonymous):

\[ 2r=3, h=4, \pi \approx 3.14 \]

OpenStudy (anonymous):

so what do i do from there solve it ?

OpenStudy (dangerousjesse):

H=4 and R=4*3/2 H=4 and R=12/2 H=4 and R=6 \[V=\pi 6 ^{2} 4\] Therefore, your answer is 452.39 meters.

OpenStudy (anonymous):

thank you, i was wondering if i had to do 3*3

OpenStudy (anonymous):

cause 2r=3

OpenStudy (dangerousjesse):

Where did you get 2r=3 from?

OpenStudy (anonymous):

wio up at the top

OpenStudy (dangerousjesse):

They probably used a different way to solve it from what I'm used to.

OpenStudy (anonymous):

Oh did you use V=1/3 (3.14) r^2 H